OpenStudy (anonymous):

what is 4 1/5 + 3 1/10?

OpenStudy (mathstudent55):

Add the whole parts together. Then add the fractions together. To add the fractions, you need a common denominator.

OpenStudy (agl202):

= 21/5 + 31/10 = ((21 × 10) + (31 × 5)) / (5 × 10) = (210 + 155) / 50 = 365/50 = 73/10 = 7 3/10

Awolflover1 (awolflover1):

the common multiple with 10 and 5 is 2 so multiply 5 x 2 and what you do to the bottom you have to do to the top so you also do 1 x 2

Awolflover1 (awolflover1):

What is 5 x 2?

OpenStudy (anonymous):

10 @Awolflover1

Awolflover1 (awolflover1):

1 x 2 is?

OpenStudy (anonymous):

2! @Awolflover1

Awolflover1 (awolflover1):